erwin's age is currently 1/5 the age of his father eight years ago the father is 13 times as old as his son was then. how old is his age now?
Answer by macston(5194)


E=Erwin's age; F= Father's age
5E=F-8
F=5E+8
F=13(E-8) Substitute for F
5E+8=13(E-8)
5E+8=13E-104 Subtract 5E from each side.
8=8E-104 Add 104 to each side.
112=8E Divide each side by 8
14=E ANSWER Erwin is 14 years old
F=13(E-8)=13(14-8)=13(6)=78 Father is 78.
CHECK
5E=F-8
5(14)=78-8
70=70